# Rethinking Debiasing: Real-World Bias Analysis and Mitigation

### v3

@@ 4.1 Debiasing with biased auxiliary model
where $(x,y)$ are samples from the training data and $CE(\cdot,\cdot)$ is the cross entropy loss. We note that the above design is introduced because it has been widely adopted by various algorithms, making it a solid entry point for our subsequent analysis. DBAM paradigm is our condensation of existing unsupervised debiasing methods, and we do not limit its scope with the above re-weighing design $W(x)$.
4.2 DBAM paradigm relies on high bias prevalence
We claim that the bias capture module of the DBAM paradigm relies on the high bias magnitude of existing benchmarks, which causes failure in real-world debiasing. It is assumed by the DBAM paradigm that the biased model $M_{b}$ predicts according to the bias within the training data, giving high loss to BC samples and low loss to BA samples [12, 9, 15, 17, 18, 20]. Existing works attribute this loss difference to the fact that spurious attributes are easier [12, 17], i.e., learn more preferentially by models, making BA the easy sample. While such a claim is true, we claim that the dominance of BA samples in the HP datasets is another vital causing factor of the loss difference, for dominant/major samples are learned more frequently than others, as shown in Figure 3(a).
However, on LP datasets, while BA samples are still easier to learn due to the biased feature, the dominant/major samples in the training data are no longer BA samples, but rather BN samples. This not only results in the loss difference between BA and BC samples decreasing but also causes low loss on BN samples, as shown in Figure 3(a). According to sample weighing scheme 4, such low loss on BN samples further leads to low weights for BN samples when training the debiased model, which is unintended as BN samples carry an abundant amount of knowledge concerning the target attribute without the interference of the spurious features. Overlooking BN samples results in server utility degradation when DBAM methods are applied to LP datasets. We empirically testify our claim in section 5.
@@ 4.3 Bias capture with feature destruction
Based on our analysis in section 4.2, we introduce a minor yet effective enhancement to the bias capture module in the existing DBAM framework. We name the refined framework as Debias in Destruction (DiD). As shown in Figure 3(b), the problem with the existing bias capture method comes from the side branch learning on BN samples of the biased auxiliary model, which not only captures the bias but also learns the target feature. This is undesired for this further causes the overlooking of BN samples when training the debiased model, as discussed in section 4.2.
@@
where $Loss_{b}$ is the original loss for training the biased model (e.g. $CE$), and $T_{fd}(\cdot)$ is the feature destruction transformation. As an example, in visual recognition tasks, the shape of objects is a basic element of human visual perception [36]. Therefore, the patch-shuffle destruction of shape [37] when capturing bias from visual recognition datasets is a feasible approach. As for NLP tasks, we adopt a word shuffling approach which we will elaborate on in Appendix D.5.
@@ 5 Experiments
In this section, with the systematic evaluation framework proposed in section 2.5, we design our experiments to answer the following questions: 1) How do existing DBAM methods and DiD perform on real-world biases? 2) Does DiD really emphasize BN samples as we intend? 3) How do debiasing methods perform on unbiased datasets? 4) How do the magnitude and prevalence of bias in datasets affect debiasing? 5) How sensitive is DiD to the hyper-parameters? (see Appendix E.2) 6) Is DiD effective on bias detection tasks as well? (see Appendix E.3) More analytical results in Appendix E.
@@ 5.1 Experimental settings
Metrics.
Following previous works [12, 9, 20], we adopt the accuracy of BC samples (BC), the average accuracy on the balanced test set (Avg), and the worst group accuracy (Worst Acc.) as the evaluation metrics. We note that DiD is not a self-contained method but rather a plug-in module for existing debiasing methods to improve from their original performances. Thus, the effectiveness of DiD should be measured as the performance gain from the base method for debiasing.
Datasets.
We adopt 8 datasets in various modalities for evaluation. Specifically, we adopt the basic setting of Colored MNIST [19] and Corrupted CIFAR10 [12] to implement the distributions within the proposed systematic evaluation framework. We also evaluated our method on more existing synthetic benchmarks who is more complex in terms of the target and spurious feature: BAR [12], NICO [15], and WaterBirds [38]. We also adopt 2 real-world NLP datasets MultiNLI [30] and CivilComments-WILDS [3], and 1 real-world image dataset CelebA [29].
Baselines.
We adopt 7 baselines, covering classic and recently proposed DBAM methods. ERM directly applies standard training on the biased datasets. LfF [12] is the first work in the DBAM paradigm. DisEnt [9] disentangles bias and intrinsic features and applies feature augmentation when training the debiased model. BE LfF and BE DisEnt were recently proposed by Lee et al. [20], and is based on LfF and DisEnt, respectively. JTT [16] is a classic DBAM method adapted to both the image and NLP domain. Group DRO [38] is a classic debiasing method with bias supervision used as an upper bound. Detailed settings can be found in Appendix D.
@@ 5.2 Main results
Evaluation on various bias distributions.
As shown in Table 1, while performing decently on HMHP distributed datasets, existing methods [12, 9, 20] fail to handle both LMLP and HMLP biases. Generally, for existing methods, the accuracy of BC samples and the average accuracy on the balanced test set is lower than the ERM baseline. This indicates the failure of exiting DBAM paradigm on the task of debiasing with low bias prevalence datasets, sacrificing utility, i.e. average accuracy, without improving worst group performance, i.e. BC accuracy. We further tested the effectiveness of DiD by combining DiD with existing DBAM methods [12, 9, 20]. As shown in Table 1, when combined with DiD, the BC and average accuracy both improve on existing HMHP benchmarks. On LMLP and HMLP datasets, the superiority of DiD is even more prominent, where BC and average accuracy both improve significantly, achieving an average of +16.6 and +11.7 for LfF and DisEnt respectively.
Table 1: The performance of our approach is presented in absolute accuracy increase of existing methods. Results show that existing DBAM methods perform poorly on LP distributions, yet our method effectively boosts the performance of existing methods across all types of biases.
Colored MNIST Corrupted CIFAR10 LMLP HMLP HMHP LMLP HMLP HMHP Algorithm BC Avg BC Avg BC Avg BC Avg BC Avg BC Avg ERM 91.1 91.7 85.2 89.8 48.5 53.4 62.5 64.3 55.9 65.1 29.4 35.4 LfF 68.4 69.7 58.0 63.3 65.6 64.6 55.0 55.4 47.7 54.1 35.3 39.0 + Ours +22.6 +21.4 +32.6 +25.8 +1.3 +3.4 +7.0 +7.3 +7.1 +8.9 +1.8 +2.5 DisEnt 73.9 74.9 66.5 72.2 68.3 67.4 55.5 56.1 52.5 54.5 36.0 39.5 + Ours +17.2 +16.5 +22.0 +16.8 +0.8 +3.1 +5.4 +5.9 +2.8 +7.1 +3.0 +3.3 BE LfF 83.6 83.5 80.0 82.3 66.9 67.6 52.1 54.0 51.0 54.0 31.5 36.6 + Ours +5.7 +6.1 +9.1 +4.9 -0.5 +0.7 +1.1 +0.2 -0.8 +0.1 +1.4 +0.8 BE DisEnt 81.1 81.0 77.6 80.2 67.5 68.5 56.6 57.2 49.1 56.3 34.2 38.6 + Ours +8.7 +9.0 +11.7 +5.5 +2.0 +2.5 +4.3 +4.2 +4.9 +5.1 +3.5 +3.2
Evaluation on complex visual features.
As shown in Table 4, our approach is not merely effective under the setting of Colored MNIST and Corrupted CIFAR10, but rather consistently demonstrating its superiority on datasets with more complex sets of target and spurious features. This shows the adaptability of DiD to more sophisticated visual data. Refer to Appendix D.2 for the metrics used.
Evaluation on real-world datasets in various modalities.
@@ 5.3 Analysis
Emphasis on BN samples.
We further validate our method by tracking the weights of samples to see if they match the purpose of our design. Figure 4(a) and 4(b) plots the average weights of all kinds of samples on HMLP distributed Colored MNIST dataset, which shows that the failure of existing methods is indeed caused by the overlooking of BN samples when training the debiased model $M_{d}$, as claimed in section 4.2. Figure 4(c) and 4(d) track the sample weight of BN samples when training LfF on HMLP distribution. As we can see, our proposed approach significantly raise the weights of the BN sample, which further demonstrates the effectiveness of our design.
(a) DisEnt
(b) LfF
(c) Colored MNIST
(d) Corrupted CIFAR10
Figure 4: Figure 4(a) and 4(b) support our claim in section 4 that existing DBAM methods tend to overlook BN samples when training on LP distributions. Figure 4(a) and 4(b) show that our approach effectively emphasizes BN samples by raising its weights.
Debias on unbiased datasets.
As we do not know how biased or is the training data biased at all in real-world scenarios, it is important to evaluate the performance of debiasing methods on unbiased training data to ensure that they do not cause severe performance degradation, if not improving the performance. As shown in Table 4, existing DBAM methods perform poorly on unbiased training data, causing severe performance degradation, yet our approach greatly boosts their performances.
@@
As shown in Figure 5, we use the correlation $Corr_{scp}$ defined in section 2 as a proxy for the bias magnitude and vary it from low to high. With the increase of the bias magnitude, the performance of LfF first increases as the data become biased, and then decreases as the bias magnitude becomes extremely high, while DiD consistently improves the performance. As shown in 5, we vary the prevalence of bias by controlling the number of biased features, which can also be viewed as an interpolation between HMLP and HMHP distribution. With the increase of the bias prevalence, the performance of LfF generally keeps increasing for its reliance on high prevalence as discussed in section 4, while DiD consistently improves the performance. Those experiments are conducted on Colored MNIST dataset.
Algorithm BAR NICO WaterBirds ERM 35.32 ±0.27 42.61 ±0.33 56.53 ±0.27 LfF 37.73 ±1.00 51.69 ±3.06 50.02 ±0.00 + Ours +3.34 ±1.69 +2.80 ±3.10 +4.47 ±0.13 DisEnt 59.11 ±1.75 39.73 ±0.58 56.75 ±4.19 + Ours +3.92 ±0.62 +16.55 ±1.29 +11.29 ±0.69 BE LfF 38.40 ±0.65 44.09 ±1.95 52.98 ±0.28 + Ours +1.08 ±1.67 +8.23 ±1.49 +1.92 ±0.12 BE DisEnt 62.74 ±1.23 39.58 ±0.91 53.85 ±2.14 + Ours +0.70 ±1.20 +13.50 ±1.62 +1.99 ±3.65
Table 4: Existing DBAM methods perform poorly on unbiased training data, while DiD greatly boosts the performance.
Algorithm Colored MNIST Corrupted CIFAR10 ERM 94.14 ± 0.21 67.91 ± 0.13 LfF 70.19 ± 1.50 52.04 ± 2.14 + Ours 93.18 ± 0.26 57.29 ± 0.22 DisEnt 75.24 ± 3.40 58.50 ± 0.20 + Ours 92.24 ± 0.44 64.58 ± 0.02 BE LfF 84.14 ± 0.61 55.64 ± 0.66 + Ours 90.02 ± 0.54 56.28 ± 0.45 BE DisEnt 80.66 ± 0.90 58.57 ± 0.12 + Ours 89.10 ± 1.28 62.97 ± 0.16
@@ 6 Conclusions and Discussion
In this work, we emphasize the importance of debias within the real world. To tackle real-world biases, we first proposed a fine-grained analysis framework to analyze dataset biases, based on which we further proposed a systematic evaluation framework for benchmarking debiasing methods under real-world biases. According to our result, we identified the insufficiency of existing methods and proposed a new approach to resolve it. Our experiments demonstrate the effectiveness of our approach. In Appendix G, we further discuss the limitations and future directions of this work.
@@ Appendix A More visualizations of biased distributions
Figure 6: Visualization of the joint distribution for datasets, where the y-axis is the target attribute and the x-axis is the spurious attribute. Figure 6(a) visualize the distribution of existing benchmarks. Figure 6(b), 6(c), 6(d), 6(e), and 6(f) visualize the distribution of real-world datasets. The biased distribution of existing benchmarks and real-world datasets is not alike.
@@ Appendix B Fine-grained evaluation framework
In this section, we elaborate on the proposed evaluation framework by mathematically and visually demonstrating the biased distribution within the biased distribution.
@@ C.1 Preliminary
Consider a classification task on binary target attribute $y^{t}\sim\{-1,+1\}$ and a binary spurious attribute $y^{s}\sim\{-1,+1\}$. Let the marginal distribution of the target and spurious attribute to be $p^{t}_{+}=P(y^{t}=+1)$ and $p^{s}_{+}=P(y^{s}=+1)$. Then the joint distribution between $y^{t}$ and $y^{s}$ can be defined according to the conditional distribution of $y^{t}$ given $y^{s}=+1$, i.e. $\tau_{+}=P(y^{t}=+1|y^{s}=+1)$. Specifically, we can derive the probability of each subgroup in the distribution:
$P(y^{s}=+1,y^{t}=+1)$ $=p^{s}_{+}\cdot\tau_{+},$ (5)
$P(y^{s}=+1,y^{t}=-1)$ $=p^{s}_{+}(1-\tau_{+}),$ (6)
$P(y^{s}=+1,y^{t}=-1)$ $=p^{t}_{+}-p^{s}_{+}\cdot\tau_{+},$ (7)
$P(y^{s}=-1,y^{t}=-1)$ $=1-p^{t}_{+}-p^{s}_{+}(1-\tau_{+})$ (8)
We assume that feature $y^{s}=+1$ and $y^{s}=-1$ is correlated $y^{t}=+1$ and $y^{t}=-1$ respectively, i.e. $\tau_{+}>p^{t}_{+}$, in the following analysis.

### v4

@@ 4.2 Existing paradigm for debiasing w/o bias label
where $(x,y)$ are samples from the training data and $CE(\cdot,\cdot)$ is the cross entropy loss. Please refer to Appendix F for detailed review on existing debiasing methods w/o bias label.
4.3 The Sparse bias capturing challenge in real-world debiasing
We claim that the effectiveness of the critical bias capture module in existing methods relies on the HP assumption of existing benchmarks, which does not generalize to LP biases in the real world. It is assumed that the biased model $M_{b}$ predicts according to the bias within the training data [19, 20], giving high loss to BC samples and low loss to BA samples [12, 9, 15, 17, 18, 23]. Existing works attribute this loss difference to the fact that spurious attributes are easier [12, 17], i.e., learn more preferentially by models, making BA the easy sample. While such a claim is true, the dominance of BA samples in the HP datasets is another vital contributing factor to the loss difference, for dominant/major samples are learned more frequently than others, as shown in Figure 3(a). The existing assumption of dense bias (BA samples are dominant) makes the bias capture process much less challenging.
However, for LP biases in the real world, while BA samples are still easier to learn due to the biased feature, the dominant/major samples in the training data are no longer BA samples, but rather BN samples. This not only results in the loss difference between BA and BC samples decreasing but also causes low loss on BN samples, as shown in Figure 3(a). According to sample weighing scheme 5, such low loss on BN samples further leads to low weights for BN samples when training the debiased model, which is unintended as BN samples carry an abundant amount of knowledge concerning the target attribute without the interference of the spurious features. In other words, the sparsity of real-world biases makes accurate bias capturing much more challenging, leading to severe degradation in the subsequent debiasing process. We empirically prove our claim in section 5.
@@ 4.4 Bias capture with feature destruction
Based on our analysis in section 4.3, we introduce a simple yet effective enhancement to the bias capture module in the existing framework. We name the refined framework as Debias in Destruction (DiD). As shown in Figure 3(b), the problem with the existing bias capture method comes from the side branch learning on BN samples of the biased auxiliary model, which not only captures the bias but also learns the target feature. This is undesired for this further causes the overlooking of BN samples when training the debiased model, as discussed in section 4.3.
@@
where $Loss_{b}$ is the original loss for training the biased model (e.g. $CE$ and $GCE$ [39]), and $T_{fd}(\cdot)$ is the feature destruction transformation. As an example, in visual recognition tasks, the shape of objects is a basic element of human visual perception [40]. Therefore, the patch-shuffle destruction of shape [41] when capturing bias from visual recognition datasets is a feasible approach. As for NLP tasks, we adopt a word shuffling approach which we will elaborate on in Appendix D.5.
@@ 5 Experiments
In this section, with the systematic evaluation framework proposed in Section 4.1, we empirically examine our findings on Sparse bias capturing challenge and the proposed approach, along with additional results on real-world debiasing.
@@ 5.1 Experimental settings
Metrics. Following previous works [12, 9, 23], we adopt the accuracy of BC samples (BC), the average accuracy on the balanced test set (Avg), and the worst group accuracy (Worst Acc.).
Datasets. We adopt 8 datasets in various modalities for evaluation. Specifically, we adopt the basic setting of Colored MNIST [21] and Corrupted CIFAR10 [12] to implement the distributions within the proposed systematic evaluation framework. We also evaluated our method on more existing synthetic benchmarks who is more complex in terms of the target and spurious feature: BAR [12], NICO [15], and WaterBirds [42]. We also adopt 2 real-world NLP datasets MultiNLI [33] and CivilComments-WILDS [3], and 1 real-world image dataset CelebA [32].
Baselines. We adopt 9 baselines, covering classic and recently proposed methods. ERM directly applies standard training on the biased datasets. LfF [12] is a pioneer work to debias w/o bias label. DisEnt [9] disentangles bias and intrinsic features and applies feature augmentation when training the debiased model. BEL, BED [23], DPR [20], DeNetDM [19] are recently proposed methods. JTT [16] is a classic method adapted to both the image and NLP domains. Group DRO [42] is a classic debiasing method with bias supervision used as an upper bound. Detailed discription in Appendix D.
@@ 5.2 Main results
Existing bias capturing degrade on LP biases, while DiD significantly boost the performance.
As shown in Table 1, while performing decently on HMHP distributed datasets, existing methods degrade on both LMLP and HMLP biases, with BC and Avg accuracy both lower than the ERM baseline. This indicates the degradation of existing bias capture module on LP biases, sacrificing utility (Avg), without improving worst group performance (BC). We further tested the effectiveness of DiD by combining DiD with existing methods. As shown in Table 1, when combined with DiD, the BC and Avg accuracy both improve on existing HMHP benchmarks. On LMLP and HMLP datasets, the superiority of DiD is even more prominent, where BC and average accuracy both improve significantly, achieving an average of +16.6 and +11.7 for LfF and DisEnt, respectively. Results on more existing methods in Appendix E.2 further show the generality of our findings.
Table 1: The performance of our approach is presented in absolute accuracy increase of existing methods. Results show that existing debiasing methods perform poorly on LP distributions, yet our method effectively boosts the performance of existing methods across all types of biases.
Colored MNIST Corrupted CIFAR10 LMLP HMLP HMHP LMLP HMLP HMHP Algorithm BC Avg BC Avg BC Avg BC Avg BC Avg BC Avg ERM 91.1 91.7 85.2 89.8 48.5 53.4 62.5 64.3 55.9 65.1 29.4 35.4 LfF 68.4 69.7 58.0 63.3 65.6 64.6 55.0 55.4 47.7 54.1 35.3 39.0 + DiD +22.6 +21.4 +32.6 +25.8 +1.3 +3.4 +7.0 +7.3 +7.1 +8.9 +1.8 +2.5 DisEnt 73.9 74.9 66.5 72.2 68.3 67.4 55.5 56.1 52.5 54.5 36.0 39.5 + DiD +17.2 +16.5 +22.0 +16.8 +0.8 +3.1 +5.4 +5.9 +2.8 +7.1 +3.0 +3.3 BEL 83.6 83.5 80.0 82.3 66.9 67.6 52.1 54.0 51.0 54.0 31.5 36.6 + DiD +5.7 +6.1 +9.1 +4.9 -0.5 +0.7 +1.1 +0.2 -0.8 +0.1 +1.4 +0.8 BED 81.1 81.0 77.6 80.2 67.5 68.5 56.6 57.2 49.1 56.3 34.2 38.6 + DiD +8.7 +9.0 +11.7 +5.5 +2.0 +2.5 +4.3 +4.2 +4.9 +5.1 +3.5 +3.2
DiD is consistently effective on complex visual features.
As shown in Table 3, our approach is not merely effective under the setting of Colored MNIST and Corrupted CIFAR10, but rather consistently effective on datasets with more complex sets of target and spurious features. This shows the adaptability of DiD to more sophisticated visual data. Refer to Appendix D.2 for the metrics used.
DiD is effective on real-world datasets in various modalities.
@@
Algorithm BAR NICO WaterBirds ERM 35.32 ±0.27 42.61 ±0.33 56.53 ±0.27 LfF 37.73 ±1.00 51.69 ±3.06 50.02 ±0.00 + DiD +3.34 ±1.69 +2.80 ±3.10 +4.47 ±0.13 DisEnt 59.11 ±1.75 39.73 ±0.58 56.75 ±4.19 + DiD +3.92 ±0.62 +16.55 ±1.29 +11.29 ±0.69 BEL 38.40 ±0.65 44.09 ±1.95 52.98 ±0.28 + DiD +1.08 ±1.67 +8.23 ±1.49 +1.92 ±0.12 BED 62.74 ±1.23 39.58 ±0.91 53.85 ±2.14 + DiD +0.70 ±1.20 +13.50 ±1.62 +1.99 ±3.65
Table 3: We experiment with three feature destruction methods with various hyperparameters on HMLP distributed dataset with LfF.
$T_{fd}$ param BC Avg N/A N/A 47.70 ± 3.58 54.15 ± 3.02 pixel-shuffle 1 51.44 ± 1.01 55.43 ± 0.20 patch-shuffle 2 51.07 ± 0.48 55.29 ± 0.27 4 49.41 ± 0.26 55.40 ± 0.26 8 54.81 ± 0.74 63.06 ± 0.77 16 49.74 ± 1.10 53.69 ± 0.31 center-occlusion 8 45.19 ± 1.41 51.61 ± 1.31 16 47.26 ± 0.54 50.94 ± 0.59 24 49.00 ± 0.80 52.60 ± 0.55 32 52.44 ± 0.87 55.76 ± 0.16
@@ 5.3 Analysis
Accuracy of bias capturing.
We further examine the accuracy of bias capturing by tracking the weights of samples to see if they align with our hypothesis in Section 4. Figure 7(a) and 7(b) plots the average weights of all kinds of samples on HMLP biases, which shows that the degradation of existing methods is indeed caused by the Sparse bias capturing challenge in section 4.3, overlooking BN samples when training the debiased model $M_{d}$. Figure 7(c) and 7(d) track the sample weight of BN samples. As we can see, DiD significantly raise the weights of the BN sample, which demonstrates more accurate bias capturing and the effectiveness of our design.
Ablation on destruction methods.
As shown in Table 3, we examine three feature destruction methods: pixel-shuffling, patch-shuffling, and center occlusion. We observed that patch-shuffle with patch-size 8 exhibits the best performance on Corrupted CIFAR10 of size 32x32.
@@
As shown in Figure 4, we use the correlation $Corr_{scp}$ defined in section 2 as a proxy for the bias magnitude and vary it from low to high. With the increase of the bias magnitude, the performance of LfF first increases as the data becomes biased, and then decreases as the bias magnitude becomes extremely high. As shown in 4, we vary the prevalence of bias by controlling the number of biased features. With the increase of the bias prevalence, the performance generally keeps increasing for its reliance on high prevalence as discussed in Section 4. In all cases DiD consistently improves the performance across the spectrum.
@@
5.4 Additional studies on real-world debiasing
We explore additional questions in real-world debiasing on the systematic evaluation framework: 1. How do debiasing methods perform on unbiased datasets? (Appendix E.3) 2. How effective is DiD on multi-bias scenarios? (Appendix E.4) 3. Is DiD effective on bias detection tasks as well? (Appendix E.5) 4. And more (Appendix E.1, E.6, E.7).
@@ 6 Conclusions and Discussion
In this work, we revisit the task of debiasing under real-world scenarios. Through solid empirical and theoretical analysis, we found a noticeable gap between existing evaluations and real-world requirements. We further fill the gap with a systematic evaluation framework for real-world debiasing. We also uncover a novel challenge in real-world debiasing, along with a simple yet effective method to address it. In Appendix G, we further discuss the limitations and future directions of this work.
@@ Appendix A More visualizations of biased distributions
Figure 5: Visualization of the joint distribution for datasets, where the y-axis is the target attribute and the x-axis is the spurious attribute. Figure 5(a) visualize the distribution of existing benchmarks. Figure 5(b), 5(c), 5(d), 5(e), and 5(f) visualize the distribution of real-world datasets. The biased distribution of existing benchmarks and real-world datasets is not alike.
@@ Appendix B Fine-grained evaluation framework
In this section, we elaborate on the proposed evaluation framework for real-world debiasing. The framework is mainly composed of three parts: Evaluation on various biased distributions in the real world, evaluation on multi-bias scenarios in the real world, and evaluation on other existing benchmarks. Covering all those aspects, we aim to provide a comprehensive and easy-to-use base for future development in the debiasing field, toward debiasing methods for real-world scenarios. Code available at https://github.com
B.1 Evaluation on various biased distributions in the real world
We mathematically and visually demonstrate the biased distribution currently included in the evaluation framework.
@@ B.1 Evaluation on various biased distributions in the real world
We note that each biased distributions are not merely the description of datasets used in this work, but rather serves as a general guide used to synthesize or sample biased datasets that reflect biases in the real world.
@@
B.2 Evaluation on multi-bias scenarios in the real world
The existence of multiple biases is another challenge in debiasing in the real world [22]. We further propose to combine multiple biases with different magnitudes and prevalence (e.g. HMLP + LMLP) together to mimick the complexity of biases in the real world. For instance, based on Corrupted CIFAR10 benchmark, containing 10 target features and 20 spurious features, we can construct a biased dataset with multiple biases of various types. Please refer to Appendix E.4 for an example of this setting.
B.3 Evaluation on other existing benchmarks
We also ensemble other popular benchmarks in the field of debiasing in an easy-to-use fashion to facilitate future research. Please refer to Appendix D.2 for details.
@@ C.1 Preliminary
Consider a classification task on the target attribute $y^{t}\sim\{a_{1}^{t},...a_{n}^{t}\}$ and a spurious attribute $y^{s}\sim\{a_{1}^{s},...a_{m}^{s}\}$. For any correlated target $a_{i}^{t}$ and spurious feature $a_{j}^{s}$, we have the marginal distribution of the target and spurious feature to be $p^{t}_{i}=P(y^{t}=a_{i}^{t})$ and $p^{s}_{j}=P(y^{s}=a_{j}^{s})$. Then the joint distribution between $y^{t}$ and $y^{s}$ can be defined according to the conditional distribution of $y^{t}$ given $y^{s}=a_{j}^{s}$, i.e. $\tau_{j}=P(y^{t}=a_{i}^{t}|y^{s}=a_{j}^{s})$. Specifically, we can derive the probability of each subgroup in the distribution:
$P(y^{s}=a_{j}^{s},y^{t}=a_{i}^{t})$ $=p^{s}_{j}\cdot\tau_{j},$ (6)
$P(y^{s}=a_{j}^{s},y^{t}\neq a_{i}^{t})$ $=p^{s}_{j}(1-\tau_{j}),$ (7)
$P(y^{s}=a_{j}^{s},y^{t}\neq a_{i}^{t})$ $=p^{t}_{i}-p^{s}_{j}\cdot\tau_{j},$ (8)
$P(y^{s}\neq a_{j}^{s},y^{t}\neq a_{i}^{t})$ $=1-p^{t}_{i}-p^{s}_{j}(1-\tau_{j})$ (9)
Furthermore, as feature $y^{s}=a_{j}^{s}$ and $y^{t}=a_{i}^{t}$ are correlated, i.e. $\tau_{j}>p^{t}_{i}$, the complement case of $y^{s}\neq a_{j}^{s}$ and $y^{t}\neq a_{i}^{t}$ is also bound to be correlated, treated as complement features.